OpenStudy (anonymous):

Find the derivative of f^(-1) for the function f at the specified value of x. (No calculator) f(x)=cosx for -pi/2≤x≤pi/2 at x=1.

OpenStudy (anonymous):

the derivative of \(\cos^{-1}(x)\) is \(-\frac{1}{\sqrt{1-x^2}}\) it is undefined at \(x=1\)